# A Survey of Progress and Challenges in Aligning LLMs with Human Intentions

### v2

@@ Abstract
Large language models (LLMs) have demonstrated exceptional capabilities in understanding and generation. However, when interacting with human instructions in real-world scenarios, LLMs still face significant challenges, particularly in accurately capturing and comprehending human instructions and intentions. This paper focuses on three challenges in LLM-based text generation tasks: instruction understanding, intention reasoning, and Reliable Dialog Generation. Regarding human complex instruction, LLMs have deficiencies in understanding long contexts and instructions in multi-round conversations. For intention reasoning, LLMs may have inconsistent command reasoning, difficulty reasoning about commands containing incorrect information, difficulty understanding user ambiguous language commands, and a weak understanding of user intention in commands. Besides, In terms of Reliable Dialog Generation, LLMs may have unstable generated content and unethical generation. To this end, we classify and analyze the performance of LLMs in challenging scenarios and conduct a comprehensive evaluation of existing solutions. Furthermore, we introduce benchmarks and categorize them based on the aforementioned three core challenges. Finally, we explore potential directions for future research to enhance the reliability and adaptability of LLMs in real-world applications.
A Survey of Progress and Challenges in Aligning LLMs
with Human Intentions
Zongyu Chang1thanks: These authors contributed equally to this work., Feihong Lu2*, Ziqin Zhu2, Qian Li1, Cheng Ji2, Zhuo Chen1, Hao Peng2, Yang Liu5, Ruifeng Xu3, Yangqiu Song4, Jianxin Li2, Shangguang Wang1 1Beijing University of Posts and Telecommunications, 2Beihang University, 3Harbin Institute of Technology, Shenzhen, 4Hong Kong University of Science and Technology, 5Chinese Academy of Sciences {changzongyu, li.qian, sgwang}@bupt.edu.cn, {lufeihong, zhuziqin, jicheng, lijx}@buaa.edu.cn, xuruifeng@hit.edu.cn, yqsong@cse.ust.hk
@@ 1 Introduction
Rapid advancements with the development of large language models (LLMs) have been experienced in the field of artificial intelligence. These models, built upon massive amounts of data and extensive computing resources, have shown impressive capabilities in understanding and generating human language. Recent advancements in LLMs, including the use of scaling laws Kaplan et al. (2020), supervised fine-tuning (SFT) Wu et al. (2021), and reinforcement learning with human feedback (RLHF) Ouyang et al. (2022), have propelled these models to new heights. Researchers have explored innovative strategies like chain-of-thought reasoning (COT) Wei et al. (2022), aiming to enhance their performance in processing and generating accurate responses. However, they still struggle with more complex interactions, especially when the input data is ambiguous, incomplete, or inconsistent. Despite improvements, issues such as content hallucination Li et al. (2023) and logical misinterpretations remain prevalent. Consequently, while LLMs show promise, they are far from flawless and require further refinement to address the challenges posed by more unpredictable and complex human instructions as follows.
@@
I. Challenge of Instruction Understanding. One of the most pressing challenges that LLMs face is instruction understanding as Figure 1(a) and Figure (2I), particularly when the user input involves complex or multi-step instructions. While models have improved in parsing relatively simple queries, they continue to encounter significant difficulties when dealing with long, context-rich instructions or when instructions are spread across multiple conversational turns. LLMs often fail to grasp subtle nuances or interpret implicit meanings embedded within the text, which leads to inaccurate or incomplete responses. Existing approaches to instruction understanding have introduced techniques like optimizing the model’s parsing abilities Teng et al. (2024), and context-aware optimization Sun et al. (2024). While these methods show promise, they often fall short when addressing the complexities and ambiguities present in instructions.
II. Challenge of Intention Reasoning. Another critical area is intention reasoning as illustrated in Figure 1(b) and Figure 2 (II), where they struggle to align the generated responses with the user’s underlying intention. Ambiguities in language, conflicting instructions, and implicit requirements often result in models producing outputs that diverge from the user’s expectations. LLMs also face difficulties when instructions are inconsistent or contain incorrect information, which challenges the model’s ability to make accurate inferences. Various strategies, including retrieval-enhanced generation and fine-tuning techniques, have been proposed to enhance reasoning capabilities, enabling the models to better handle inconsistent or incomplete instructions. However, these methods often introduce new challenges related to bias and the inability to fully resolve conflicts in user input, further complicating the alignment between generated content and user expectations.
III. Challenge of Reliable Dialog Generation. The final major challenge is the reliable dialog generation, which pertains to the accuracy, ethical considerations, and stability of the content they produce, such as Figure 1(c) and Figure 2 (III. While LLMs are generally capable of generating coherent and contextually relevant outputs, they sometimes exhibit instability, generating content that is factually incorrect, logically inconsistent, or ethically questionable. This challenge is exacerbated by the model’s inability to recognize uncertainty, which can lead to overconfident but inaccurate outputs. Recent efforts to address this issue involve techniques like uncertainty-aware fine-tuning and using external tools to evaluate output credibility. However, these approaches struggle to provide a comprehensive and reliable solution, especially in complex or dynamic contexts.
@@
forked edges, for tree= grow=east, reversed=true, base=left, font=, rectangle, draw=hidden-draw, rounded corners, align=left, minimum width=4em, edge+=darkgray, line width=1pt, inner xsep=2pt, inner ysep=3pt, ver/.style=rotate=90, child anchor=north, anchor=center , [ Gap Between LLMs and Human Intention, ver, color=hidden-draw, fill=mygray!30, text=black [ Instruction
Understanding (§2), color=myblue!100, fill=myblue!80, text width=6em, text=black [ Long-text
Comprehension (§2.1), color=bluelight1!60, fill=bluelight1!60, text width=7em, text=black [ Information
Focusing, color=bluelight1!40, fill=bluelight1!40, text width=4em, text=black [ 1) Attention Sparsification Beltagy et al. (2020),
2) Attention Optimization Chen et al. (2024),
3) Position independent Training He et al. (2024a) , understanding_leaf, text width=26.5em ] ] [ Multipath
Optimization, color=bluelight1!40, fill=bluelight1!40, text width=4em, text=black [ 1) SFT and RL Zhang et al. (2024b)
2) Retrieval-Augmented Li et al. (2024b)
3) Recurrent Sequence Optimization Gu and Dao (2023) , understanding_leaf, text width=26.5em ] ] ] [ Multi-turn Conversation
Handling (§2.2), color=bluelight1!60, fill=bluelight1!60, text width=7em, text=black [ Soft
Fine-tuning, color=bluelight1!40, fill=bluelight1!40, text width=4em, text=black [ Multi-turn SFT Teng et al. (2024), Parrot Sun et al. (2024) , understanding_leaf, text width=26.5em ] ] [ Reforcement
Learning, color=bluelight1!40, fill=bluelight1!40, text width=4em, text=black [ ArCHer Zhou et al. (2024) , understanding_leaf, text width=26.5em ] ] ] ] [ Intention
Reasoning (§3), color=mygreen!80, fill=mygreen!80, text width=6em, text=black [ Inconsistent Instruction
Resolution (§3.1), color=mygreen!60, fill=mygreen!60, text width=7em, text=black [ Knowledge
Updating, color=mygreen!40, fill=mygreen!40, text width=4em, text=black [ SituatedQA Zhang and Choi (2021), CDConv Zheng et al. (2022) , reasoing_leaf, text width=26.5em ] ] [ Confidence
Calibration, color=mygreen!40, fill=mygreen!40, text width=4em, text=black [ CD2 Jin et al. (2024), MacNoise Hong et al. (2024) , reasoing_leaf, text width=26.5em ] ] ] [ Misinformation
Reasoning (§3.2), color=mygreen!60, fill=mygreen!60, text width=7em, text=black [ Targeted
Fine-tuning, color=mygreen!40, fill=mygreen!40, text width=4em, text=black [ CKL Jang et al. (2022), StreamingQA Liska et al. (2022), BIPIA Yi et al. (2025),
KC-LLMs Wang et al. (2023), CAR Weller et al. (2024) , reasoing_leaf, text width=26.5em ] ] ] [ Fuzzy Language
Interpretation (§3.3), color=mygreen!60, fill=mygreen!60, text width=7em, text=black [ Clue
Engineering , color=mygreen!40, fill=mygreen!40, text width=4em, text=black [ Folkscope Yu et al. (2023), Miko Lu et al. (2024), ATC Deng et al. (2023) , reasoing_leaf, text width=26.5em ] ] ] [ Intention
Clarification
Failure (§3.4) , color=mygreen!60, fill=mygreen!60, text width=7em, text=black [ Deep
Reasoning , color=mygreen!40, fill=mygreen!40, text width=4em, text=black [ DeepSeek-R1 Guo et al. (2025), S1 Muennighoff et al. (2025),
SoulChat Chen et al. (2023), MoChat Mo et al. (2024) , reasoing_leaf, text width=26.5em ] ] ] ] [ Reliable
Generation (§4), color=lightkhaki!80, fill=lightkhaki!80, text width=6em, text=black [ Response Stability (§4.1), color=lightkhaki!60, fill=lightkhaki!60, text width=7em, text=black [ Fine-tuning
LLMs, color=lightkhaki!40, fill=lightkhaki!40, text width=4em, text=black [ EDL Sensoy et al. (2018), DER Amini et al. (2020),
ConformalFactuality Mohri and Hashimoto (2024), LUQ Zhang et al. (2024a) , reliable_leaf, text width=26.5em ] ] [ External Tools, color=lightkhaki!40, fill=lightkhaki!40, text width=4em, text=black [ SUE Liu et al. (2024), CalibrateMath Lin et al. , reliable_leaf, text width=26.5em ] ] ] [ Alignment (§4.2), color=lightkhaki!60, fill=lightkhaki!60, text width=7em, text=black [ Data Cleaning
and Curation, color=lightkhaki!40, fill=lightkhaki!40, text width=4em, text=black [ Stochastic Parrots Bender et al. (2021) , DeepSoftDebias Rakshit et al. , reliable_leaf, text width=26.5em ] ] [ RL-based
Alignment , color=lightkhaki!40, fill=lightkhaki!40, text width=4em, text=black [ PPO Ouyang et al. (2022), DPO Zeng et al. (2024), GRPO Shao et al. (2024),
RLAIF Lee et al. (2024) , reliable_leaf, text width=26.5em ] ] [ In-context
Alignment, color=lightkhaki!40, fill=lightkhaki!40, text width=4em, text=black [ URIAL Lin et al. (2024), ICA Huang et al. (2024a) , reliable_leaf, text width=26.5em ] ] ] ] ]
@@
Present Survey.
Facing these challenges, there is an increasing need for focused research on LLMs and their interaction with human instructions and intentions. This paper systematically analyzes LLMs’ performance in processing human instructions, highlighting three key areas: user instruction understanding, intention comprehension and reasoning, and reliable dialog generation. While existing review papers address model training, fine-tuning, and specific aspects of LLMs’ capabilities Lou et al. (2024); Plaat et al. (2024); Huang et al. (2024b), our focus is on the LLMs’ ability to understand and reason about user intentions. Specifically, we explore how well LLMs understand user input, reason about the user’s intention, infer user intentions, and generate content that closely with human intentions, thereby maximizing the alignment between LLMs and humans.
Comparison with Previous Surveys.
While the gap between human intention and LLMs is a core challenge in generative AI, many studies focus on specific aspects of the issue, lacking a comprehensive overview. These works offer valuable insights but do not provide a systematic summary of the field. Lou et al. Lou et al. (2024) primarily address instruction following challenges in LLMs without delving into the reasoning capabilities for complex user instructions. Gao et al. analyze the four stages of human-machine LLM interaction (planning, facilitation, iteration, and testing) but overlook LLM’s understanding of user instructions. Xu et al. Xu et al. examine the impact of various memory conflicts on LLM-generated content credibility and performance, yet do not consider reasoning or intention comprehension. Plaat et al. Plaat et al. (2024) focus on LLM Reasoning for basic mathematical problems, without exploring its applicability to broader fields. Shorinwa et al. Shorinwa et al. (2024) provide an initial analysis of LLMs uncertainty quantification, but exclude user input instructions. In contrast, our survey offers a more comprehensive perspective, as shown in Figure 2 and Figure 3, with a unique classification and systematic analysis of instruction processing, while addressing current solutions to key challenges.
Survey Organization.
@@ 2.1 Long-Text Comprehension
Information Focusing.
Improving LLM’s ability to focus on important information in long texts involves several methods: 1) Sparsifying attention to concentrate on critical information Beltagy et al. (2020). 2) Optimizing attention to minimize redundancy and emphasize core content Chen et al. (2024). 3) Training with location-independent tasks to enhance the ability to search and react to relevant information in long contexts He et al. (2024a).
Multipath Optimization.
Various methods can enhance LLMs on long-context tasks: 1) Pre-training with extended context windows and reinforcement learning for fine-tuning to optimize long-context understanding Zhang et al. (2024b). 2) Combining retrieval-based models with generative models on long-context tasks Li et al. (2024b). 3) Leveraging cyclic sequence models’ linear scaling property for better inference efficiency Gu and Dao (2023).
@@ 2.2 Multi-Turn Conversation Handling
Multi-turn conversation serves as a fundamental interaction mode between LLMs and humans. Given the challenges users face in providing complete and precise instructions in a single turn, they often opt to refine and clarify their intentions incrementally through iterative exchanges. However, due to the characteristics of real-world conversations, such as the constant changes in user intentions and long-distance dependencies, LLMs still faces significant challenges in achieving coordinated multi-round interactions with humans.This paper categorizes the challenges faced by existing LLMs when understanding multi-turn conversations into three categories, as follows: 1) Capability Weakening. Current supervised instruction fine-tuning (SIFT) and RLHF may even impair multi-turn capabilitiesWang et al. (2024), with models struggling on complex reasoning tasks that span multiple rounds, such as those requiring evidence collection and conclusions Banatt et al. (2024). Additionally, multi-turn dialogs increase the vulnerability of LLMs to adversarial attacks, where malicious users can mask harmful intentions across multiple rounds, leading to the generation of misleading or harmful content Agarwal et al. (2024). 2) Error Propagation. Instruction comprehension errors accumulate across rounds, leading to an escalating failure rate in subsequent responses He et al. (2024b), which may snowball into larger issues such as biased or incorrect outputs Fan et al. (2024). 3) Incorrect Relevance Judgment (Figure 5 Q.1) LLMs often struggle to identify relevant content in multi-turn dialogs, failing to properly link content from previous rounds or to discern ellipsis and implicit meaning inherent in user commands Sun et al. (2024).
To solve above challenges, this paper categorizes existing solutions into two types: supervised fine-tuning methods using multi-turn dialogue data, enhanced by techniques like optimized instruction parsing Teng et al. (2024) and context-aware preference strategies Sun et al. (2024); and reinforcement learning methods tailored for multi-turn dialogue, with improvements such as hierarchical reinforcement learning Zhou et al. (2024).
@@ 3.1 Inconsistent Instruction Reasoning
In natural language communication, humans easily identify inconsistencies using context and prior knowledge, whereas LLMs struggle, often accept contradictory inputs, and generate unreliable answers. This phenomenon has been observed across multiple question-answering generation tasks Li et al. (2024a); Zheng et al. (2022), and we categorize the causes of this problem according to the scenarios in which it occurs as follows: 1) Ignoring input errors (Figure 6 Q.1). The model ignores the input errors and gives an answer, resulting in the model assigning the same weight to each context given by the user, which in turn affects the generation of the answer. 2) Inability to detect user inconsistencies (Figure 6 Q.2). In the premise that the model has learned the knowledge, the model still has difficulty detecting user inconsistencies. To address inconsistent instruction reasoning issues, existing solutions primarily adopt the following two approaches:
Knowledge Updating.
SituatedQA Zhang and Choi (2021) attempts to enhance model performance by updating the knowledge base. Additionally, CDConv Zheng et al. (2022) simulates common user behaviors to trigger chatbots through an automated dialogue generation method, generating contradictions for training purposes.
Confidence Calibration.
Given the high cost of data annotation and model fine-tuning, some researchers have sought alternative approaches by introducing additional processing techniques. CD2 Jin et al. (2024) maximizes probabilistic output and calibrates model confidence under knowledge conflicts using conflict decoupling and comparison decoding methods.

### v3

# A Survey of LLM Alignment: Instruction Understanding, Intention Reasoning, and Reliable Generation

@@ Abstract
Large language models have demonstrated exceptional capabilities in understanding and generation. However, in real-world scenarios, users’ natural language expressions are often inherently fuzzy, ambiguous, and uncertain, leading to challenges such as vagueness, polysemy, and contextual ambiguity. This paper focuses on three challenges in LLM-based text generation tasks: instruction understanding, intention reasoning, and reliable dialog generation. Regarding human complex instruction, LLMs have deficiencies in understanding long contexts and instructions in multi-round conversations. For intention reasoning, LLMs may have inconsistent command reasoning, difficulty reasoning about commands containing incorrect information, difficulty understanding user ambiguous language commands, and a weak understanding of user intention in commands. Besides, In terms of Reliable Dialog Generation, LLMs may have unstable generated content and unethical generation. To this end, we classify and analyze the performance of LLMs in challenging scenarios and conduct a comprehensive evaluation of existing solutions. Furthermore, we introduce benchmarks and categorize them based on the aforementioned three core challenges. Finally, we explore potential directions for future research to enhance the reliability and adaptability of LLMs in real-world applications.
keywords:
Instruction Understanding, Intention Reasoning, Reliable Dialog Generation
11footnotetext: * These authors contributed equally to this work.
[…]
@@ 1 Introduction
Rapid advancements with the development of large language models (LLMs) have been experienced in the field of artificial intelligence. These models, built upon massive amounts of data and extensive computing resources, have shown impressive capabilities in understanding and generating human language. Recent advancements in LLMs, including the use of scaling laws [1], supervised fine-tuning (SFT) [2], and reinforcement learning with human feedback (RLHF) [3], have propelled these models to new heights. Researchers have explored innovative strategies like chain-of-thought reasoning (COT) [4], aiming to enhance their performance in processing and generating accurate responses. However, they still have struggled when interacting with human instructions in real-world scenarios due to the fuzzy, uncertain, and complex nature of human instructions, especially when the input data is fuzzy, incomplete, or inconsistent. Despite improvements, issues such as content hallucination [5] and logical misinterpretations remain prevalent. Consequently, while LLMs show promise, they are far from flawless and require further refinement to address the challenges posed by more unpredictable and complex human instructions as follows.
@@
1.1 Challenge of Instruction Understanding
One of the most pressing challenges that LLMs face is instruction understanding as Figure 1(a) and Figure (2I), particularly when the user input involves complex or multi-step instructions. While models have improved in parsing relatively simple queries, they continue to encounter significant difficulties when dealing with long, context-rich instructions or when instructions are spread across multiple conversational turns. LLMs often fail to grasp subtle nuances or interpret implicit meanings embedded within the text, which leads to inaccurate or incomplete responses. Existing approaches to instruction understanding have introduced techniques like optimizing the model’s parsing abilities [6], and context-aware optimization [7]. While these methods show promise, they often fall short when addressing the complexities and ambiguities present in instructions.
1.2 Challenge of Intention Reasoning
Another critical area is intention reasoning as illustrated in Figure 1(b) and Figure 2 (II), where they struggle to align the generated responses with the user’s underlying intention. Ambiguities in language, conflicting instructions, and implicit requirements often result in models producing outputs that diverge from the user’s expectations. LLMs also face difficulties when instructions are inconsistent or contain incorrect information, which challenges the model’s ability to make accurate inferences. Various strategies, including retrieval-enhanced generation and fine-tuning techniques, have been proposed to enhance reasoning capabilities, enabling the models to better handle inconsistent or incomplete instructions. However, these methods often introduce new challenges related to bias and the inability to fully resolve conflicts in user input, further complicating the alignment between generated content and user expectations.
1.3 Challenge of Reliable Dialog Generation.
The final major challenge is the reliable dialog generation, which pertains to the accuracy, ethical considerations, and stability of the content they produce, such as Figure 1(c) and Figure 2 (III. While LLMs are generally capable of generating coherent and contextually relevant outputs, they sometimes exhibit instability, generating content that is factually incorrect, logically inconsistent, or ethically questionable. This challenge is exacerbated by the model’s inability to recognize uncertainty, which can lead to overconfident but inaccurate outputs. Recent efforts to address this issueinvolve techniques like uncertainty-aware fine-tuning and using external tools to evaluate output credibility. However, these approaches struggle to provide a comprehensive and reliable solution, especially in complex or dynamic contexts.
@@ 1.2 Challenge of Intention Reasoning
Present Survey
Facing these challenges, there is an increasing need for focused research on LLMs and their interaction with human instructions and intentions. This paper systematically analyzes LLMs’ performance in processing human instructions, highlighting three key areas: user instruction understanding, intention comprehension and reasoning, and reliable dialog generation. While existing review papers address model training, fine-tuning, and specific aspects of LLMs’ capabilities [92, 93, 94], our focus is on the LLMs’ ability to understand and reason about user intentions. Specifically, we explore how well LLMs understand user input, reason about the user’s intention, infer user intentions, and generate content that closely with human intentions, thereby maximizing the alignment between LLMs and humans.
Comparison with Previous Surveys
While the gap between human intention and LLMs is a core challenge in generative AI, many studies focus on specific aspects of the issue, lacking a comprehensive overview. These works offer valuable insights but do not provide a systematic summary of the field. Lou et al. [92] primarily address instruction following challenges in LLMs without delving into the reasoning capabilities for complex user instructions. Xu et al. [95] examine the impact of various memory conflicts on LLM-generated content credibility and performance, yet do not consider reasoning or intention comprehension. Plaat et al. [93] focus on LLM Reasoning for basic mathematical problems, without exploring its applicability to broader fields. Shorinwa et al. [96] provide an initial analysis of LLMs uncertainty quantification, but exclude user input instructions. In contrast, our survey offers a more comprehensive perspective, as shown in Figure 2 and Figure 3, with a unique classification and systematic analysis of instruction processing, while addressing current solutions to key challenges.
Survey Organization
@@ 2.1 Long-Text Comprehension
Information Focusing
Improving LLM’s ability to focus on important information in long texts involves several methods: 1) Sparsifying attention to concentrate on critical information [8]. 2) Dynamically adjusting attention based on the current task [9]. 3) Optimizing attention to minimize redundancy and emphasize core content [10]. 4) Training with location-independent tasks to enhance the ability to search and react to relevant information in long contexts [11].
Multipath Optimization
Various methods can enhance LLMs on long-context tasks: 1) Pre-training with extended context windows and reinforcement learning for fine-tuning to optimize long-context understanding [12, 13, 14, 15]. 2) Combining retrieval-based models with generative models on long-context tasks [16, 17]. 3) Leveraging cyclic sequence models’ linear scaling property for better inference efficiency [18]. 4) Using external memory to store and retrieve long-context information [19]. 5) Mimicking brain memory hierarchies to improve long-context processing efficiency [20].
@@ 2.2 Multi-Turn Conversation Handling
Multi-turn conversation serves as a fundamental interaction mode between LLMs and humans. Given the challenges users face in providing complete and precise instructions in a single turn, they often opt to refine and clarify their intentions incrementally through iterative exchanges. However, due to the characteristics of real-world conversations, such as the constant changes in user intentions and long-distance dependencies, LLMs still faces significant challenges in achieving coordinated multi-round interactions with humans.This paper categorizes the challenges faced by existing LLMs when understanding multi-turn conversations into three categories, as follows: 1) Capability Weakening. Current supervised instruction fine-tuning (SIFT) and RLHF may even impair multi-turn capabilities[97], with models struggling on complex reasoning tasks that span multiple rounds, such as those requiring evidence collection and conclusions [98]. Additionally, multi-turn dialogs increase the vulnerability of LLMs to adversarial attacks, where malicious users can mask harmful intentions across multiple rounds, leading to the generation of misleading or harmful content [99]. 2) Error Propagation. Instruction comprehension errors accumulate across rounds, leading to an escalating failure rate in subsequent responses [100], which may snowball into larger issues such as biased or incorrect outputs [101]. 3) Incorrect Relevance Judgment (Figure 5 Q.1) LLMs often struggle to identify relevant content in multi-turn dialogs, failing to properly link content from previous rounds or to discern ellipsis and implicit meaning inherent in user commands [7].
To solve above challenges, this paper categorizes existing solutions into two types: supervised fine-tuning methods using multi-turn dialogue data, enhanced by techniques like optimized instruction parsing [6] and context-aware preference strategies [7]; and reinforcement learning methods tailored for multi-turn dialogue, with improvements such as hierarchical reinforcement learning [28].
@@ 3.1 Inconsistent Instruction Reasoning
In natural language communication, humans easily identify inconsistencies using context and prior knowledge, whereas LLMs struggle, often accept contradictory inputs, and generate unreliable answers. This phenomenon has been observed across multiple question-answering generation tasks [33, 30], and we categorize the causes of this problem according to the scenarios in which it occurs as follows: 1) Ignoring input errors (Figure 6 Q.1). The model ignores the input errors and gives an answer, resulting in the model assigning the same weight to each context given by the user, which in turn affects the generation of the answer. 2) Inability to detect user inconsistencies (Figure 6 Q.2). In the premise that the model has learned the knowledge, the model still has difficulty detecting user inconsistencies. To address inconsistent instruction reasoning issues, existing solutions primarily adopt the following two approaches:
Knowledge Updating
SituatedQA [29] attempts to enhance model performance by updating the knowledge base. ContraDoc [33] trains models to identify contradictions in long texts using human-annotated datasets. Additionally, CDConv [30] simulates common user behaviors to trigger chatbots through an automated dialogue generation method, generating contradictions for training purposes.See Figure 3 for more methods.
Confidence Calibration
Given the high cost of data annotation and model fine-tuning, some researchers have sought alternative approaches by introducing additional processing techniques. CD2 [34] maximizes probabilistic output and calibrates model confidence under knowledge conflicts using conflict decoupling and comparison decoding methods. MacNoise [39] enhances contradiction retrieval in an augmented generative system by explicitly fine-tuning a discriminator or prompting LLMs to improve contradiction discrimination. See Figure 3 for more methods.